# Lorentzian Residual Neural Networks

### v1

@@ Abstract.
Residual Neural Networks, Hyperbolic Geometry, Deep Neural Networks, Foundation Model
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Machine learningccs: Computing methodologies Knowledge representation and reasoningccs: Mathematics of computing Geometric topology
@@ 5.2. Adaptation to Graph Transformers
Method Chameleon Squirrel Actor
@@
Method CIFAR-10 (ACC) CIFAR-100 (ACC)
@@ Appendix A Proof of Theoretical Results
$c_{v}=\frac{\langle\mathbf{x},\mathbf{u}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{x}\rangle_{\mathcal{L}}}=\frac{\left[-2\beta a_{1}^{2}-\beta a_{n+1}^{2}+\displaystyle\sum_{i=2}^{n}\beta a_{i}^{2}\right]^{T}}{-1/K-\sqrt{-1/K}a_{1}}$
@@
$c_{v}^{\prime}=\frac{\langle\mathbf{y},\mathbf{u}^{\prime}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{y}\rangle_{\mathcal{L}}}=\frac{\left[-2\beta a_{1}^{2}-\beta a_{n+1}^{2}+\displaystyle\sum_{i=2}^{n}\beta a_{i}^{2}\right]^{T}}{-1/K-\sqrt{-1/K}a_{1}}.$
@@
    For parallel transport method: The isometry from Lorentz model to Klein model is given by the map
    $\varphi_{K}(\mathbf{x})=\mathbf{x}_{t}/x_{s}$
    . Thus given the fact that geodesics in the Klein model are Euclidean straight lines, it suffices to show that $\mathbf{z}_{s}=\lambda\mathbf{m}_{s}$ for some $\lambda\in\mathbb{R}$. Now we can compute:
@@
    where $c_{u}=\frac{\cosh^{-1}\left(y_{t}\sqrt{-K}\right)}{\sqrt{-y_{t}^{2}K-1}}$, $c_{v}=\frac{\langle\mathbf{x},\mathbf{u}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{x}\rangle_{\mathcal{L}}},\mathbf{y}^{\prime}=\mathbf{y}+y_{t}\sqrt{-K}\mathbf{o},\mathbf{x}^{\prime}=\mathbf{o}+\mathbf{x},\alpha=\sqrt{-K}||\mathbf{v}||_{\mathcal{L}}$. Clearly $\mathbf{y}^{\prime}_{s}=\mathbf{y}_{s},\mathbf{x}^{\prime}_{s}=\mathbf{x}_{s}$. So $w_{x}=\cosh(\alpha)+c_{v}$ and $w_{y}=\frac{\sinh(\alpha)}{\alpha}c_{u}$ gives the desired result.
@@
    For space addition method: This follows immediately from the isometry to Klein model.
@@ B.1. Graph datasets
B.2. Image datasets
For the image classification tasks, we used the splits implemented in PyTorch for CIFAR-10 and CIFAR-100, where there are 50,000 training images and 10,000 testing images. For OOD-dection, we followed the set up in (van Spengler et al., 2023), where we use the same data splits as in (van Spengler et al., 2023) and (Liu et al., 2020).
@@ C.1.1. GNN hyperparameters
For heteophilic graph datasets, we used a curvature of $-1$ and performed a grid search in the following search space: dimension within $\{16,32,64\}$; learning rate within $\{0.001,0.01\}$; dropout rate within $\{0,0.1,0.2,0.3\}$; weight-decay rate within $\{0,1e-4,1e-3\}$; number of layers within $\{3,4,5,6,7,8\}$. While many values were searched, we find that the performance of LResNet and the baselines depend mostly on the performance of the base HyboNet without residual connection. As a result, we used the same hyperparameters as in (Chen et al., 2021).
@@ C.1.2. Hyperbolic SGFoermer Hyperparameters
We performed a grid search in the following search space with a constant curvature of -1.0:
  • dimensions within $\{32,64\}$ for Chameleon and Squirrel, fixed dimension of 96 for Film
  • learning rate within $\{0.001,0.01\}$
  • weight-decay rate within $\{0,1e-4,1e-3\}$
  • dropout rate within $\{0.3,0.4,0.5,0.6\}$
  • number of layers within $\{2,3,4,5,6,7,8\}$
@@ C.2. Implementation details for image datasets
For hyperparameters in the classification task, we performed a grid search where learning rate is within the set $\{0.1,0.01,0.001\}$ and weight decay is within the set $\{0,5e-4,5e-3\}$. For OOD-detection, we trained the model with the same hyperparameter as in (van Spengler et al., 2023) and using the Riemmanian Adam optimizer as in (van Spengler et al., 2023) as well. In both cases, we performed search for curvature in the set $\{0.1,1,1.5\}$ and the scaling coefficient in Equation 10 in the set $\{0.5,1.0,2.0,\mathrm{trainable}\}$.

### v2

@@ Abstract.
Residual Neural Networks; Hyperbolic Geometry; Deep Neural Networks; Foundation Model
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709292isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Machine learningccs: Computing methodologies Knowledge representation and reasoningccs: Mathematics of computing Geometric topology
@@ 5.2. Adaptation to Graph Transformers
Dataset Chameleon Squirrel Actor
@@
Dataset CIFAR-10 (ACC) CIFAR-100 (ACC)
@@ Appendix A Proof of Theoretical Results
Proof of Lemma 4.1
Proof.
Note that
$-K||w_{x}\mathbf{x}+w_{y}\mathbf{y}||_{\mathcal{L}}^{2}$
$=-K\left(-(w_{x}x_{t}+w_{y}y_{t})^{2}+||w_{x}\mathbf{x}_{s}+w_{y}\mathbf{y}_{s}||^{2}\right)$
$=K||w_{x}\mathbf{x}||_{\mathcal{L}}^{2}+K||w_{y}\mathbf{y}||_{\mathcal{L}}^{2}-2K\langle w_{x}\mathbf{x},w_{y}\mathbf{y}\rangle_{\mathcal{L}}$
$=w_{x}^{2}+w_{y}^{2}-2K\Gamma$
$>w_{x}^{2}+w_{y}^{2}-2K(w_{x}w_{y}||\mathbf{x}_{s}||||\mathbf{y}_{s}||-w_{x}w_{y}\langle\mathbf{x}_{s}.\mathbf{y}_{s}\rangle)$
(Cauchy Schwarz Inequality) $\geq w_{x}^{2}+w_{y}^{2},$
where
(13) $\begin{split}\Gamma&=w_{x}\sqrt{\|\mathbf{x}_{s}\|^{2}-\frac{1}{K}}\cdot w_{y}\sqrt{||\mathbf{y}_{s}||^{2}-\frac{1}{K}}-w_{x}w_{x}\langle\mathbf{x}_{s},\mathbf{y}_{s}\rangle\end{split}$
By taking the square roots we have the desired inequality. ∎
@@
$c_{v}$ $=\frac{\langle\mathbf{x},\mathbf{u}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{x}\rangle_{\mathcal{L}}}$
$=\frac{\left[-2\beta a_{1}^{2}-\beta a_{n+1}^{2}+\displaystyle\sum_{i=2}^{n}\beta a_{i}^{2}\right]^{T}}{-1/K-\sqrt{-1/K}a_{1}}$
@@
$c_{v}^{\prime}$ $=\frac{\langle\mathbf{y},\mathbf{u}^{\prime}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{y}\rangle_{\mathcal{L}}}$
$=\frac{\left[-2\beta a_{1}^{2}-\beta a_{n+1}^{2}+\displaystyle\sum_{i=2}^{n}\beta a_{i}^{2}\right]^{T}}{-1/K-\sqrt{-1/K}a_{1}}.$
@@
    For parallel transport method: First note that the isometry from Lorentz model to Klein model is given by the map
    $\varphi_{K}(\mathbf{x})=\mathbf{x}_{t}/x_{s}.$
    Thus given the fact that geodesics in the Klein model are Euclidean straight lines, it suffices to show that $\mathbf{z}_{s}=\lambda\mathbf{m}_{s}$ for some $\lambda\in\mathbb{R}$. Then we can compute:
@@
    where
    $c_{u}=\frac{\cosh^{-1}\left(y_{t}\sqrt{-K}\right)}{\sqrt{-y_{t}^{2}K-1}},c_{v}=\frac{\langle\mathbf{x},\mathbf{u}\rangle_{\mathcal{L}}}{-1/K-\langle\mathbf{o},\mathbf{x}\rangle_{\mathcal{L}}},$
    $\mathbf{y}^{\prime}=\mathbf{y}+y_{t}\sqrt{-K}\mathbf{o},\mathbf{x}^{\prime}=\mathbf{o}+\mathbf{x},\alpha=\sqrt{-K}||\mathbf{v}||_{\mathcal{L}}.$
    Clearly $\mathbf{y}^{\prime}_{s}=\mathbf{y}_{s},\mathbf{x}^{\prime}_{s}=\mathbf{x}_{s}$. So $w_{x}=\cosh(\alpha)+c_{v}$ and $w_{y}=\frac{\sinh(\alpha)}{\alpha}c_{u}$ gives the desired result.
@@
    For space addition method: This follows immediately from the isometry to the Klein model.
@@ Appendix B Dataset Processing Details
Image datasets
For the image classification tasks, we used the splits implemented in PyTorch, where there are 50,000 training images and 10,000 testing images. For OOD-dection, we followed the set up in (van Spengler et al., 2023), where we use the same data splits as in (van Spengler et al., 2023) and (Liu et al., 2020).
@@ C.1.1. GNN hyperparameters
For heteophilic graph datasets, we used a curvature of $-1$ and performed a grid search in the following search space: dimension within $\{16,32,64\}$; learning rate within $\{0.001,0.01\}$;dropout rate within $\{0,0.1,0.2,0.3\}$; weight-decay rate within $\{0,1e-4,1e-3\}$; number of layers within $\{3,4,5,6,7,8\}$. While many values were searched, we find that the performance of LResNet and the baselines depend mostly on the that of the base HyboNet without residual connection. As a result, we used the same hyperparameters as in (Chen et al., 2021).
@@ C.1.2. Hyperbolic SGFoermer Hyperparameters
We performed a grid search with a constant curvature of -1.0: dimensions within $\{32,64\}$ for Chameleon and Squirrel, fixed dimension of 96 for Film; learning rate within $\{0.001,0.01\}$; weight-decay rate within $\{0,1e-4,1e-3\}$; dropout rate within $\{0.3,0.4,0.5,0.6\}$; number of layers within $\{2,3,4,5,6,7,8\}$.
@@ C.2. Implementation details for image datasets
We performed a grid search where learning rate is within the set $\{0.1,0.01,0.001\}$ and weight decay is within the set $\{0,5e-4,5e-3\}$. For OOD-detection, we trained the model with the same hyperparameter as in (van Spengler et al., 2023) and using the Riemmanian Adam optimizer as in (van Spengler et al., 2023) as well. In both cases, we performed search for curvature in the set $\{0.1,1,1.5\}$ and the scaling coefficient in Equation 10 in the set $\{0.5,1.0,2.0,\mathrm{trainable}\}$.